OpenStudy (anonymous):

Choose the Standard Form of the equation y = -one over four x - 6. Answer one over four x + y = -6 4x + y = -6 x + 4y = -24 -4x + y = -24

OpenStudy (anonymous):

please explain each step not the answer

Directrix (directrix):

x + 4y = -24 ========= y = (-1/4) [ x] - 6 multiply each member of the equation by 4 to clear fractions 4y = -1x - 24 Add -1x to both sides of the equation 4y + 1x = -24 x + 4y = -24 Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.